STRUCTURE OF THE SEVOFLURANE-BENZENE COMPLEX AS DETERMINED BY CHIRPED-PULSE FTMW SPECTROSCOPY

\underline{NATHAN A. SEIFERT}, DANIEL P. ZALESKI, JUSTIN L. NEILL, BROOKS H. PATE, Department of Chemistry, University of Virginia, McCormick Rd., Charlottesville, VA 22904-4319; ALBERTO LESARRI, MONTSERRAT VALLEJO, Departamento de Química Física y Química Inorgánica, Facultad de Ciencias, Universidad de Valladolid, E-47011 Valladolid, Spain; EMILIO J. COCINERO, FERNANDO CASTANO, Departamento de Química Física, Facultad de Ciencia y Tecnología, Universidad del País Vasco (UPV/EHU), Campus de Leioa, Ap. 644, E-48080 Bilbao, Spain.

Following previous microwave studies on sevoflurane monomer by Suenram et al.^{a} and Vega-Toribio et al.^{b} we report the broadband rotational spectrum of sevoflurane clustered with benzene. The structure assigned is consistent with a C-H⋯ π interaction between the benzene ring and the (CF_{3})_{2}C-H hydrogen on sevoflurane. The spectrum of this species is complicated by the six-fold internal rotation of the benzene ring over the C_{1} framework of sevoflurane. The six-fold tunneling falls into a high effective barrier case where there are several bound torsional levels. The tunneling spectrum has been successfully analyzed using the BELGI internal rotation program and a barrier to internal rotation of the benzene against sevoflurane of 32.5 cm^{−1} has been determined. Structural information about the complex has been obtained by studying the complex of sevoflurane with benzene-d_{1}. For this complex, six unique isomers are observed making it possible to determine the positions of the benzene H-atoms in the complex. Combination of these hydrogen r_{s} positions with the sevoflurane monomer r_{s} coordinates reported by Lesarri et al.^{c} results in a substitution structure in excellent agreement with the ab initio results. Finally, initial microwave results on two sevoflurane dimer species will also be presented.

aR. D. Suenram, D. J. Brugh, F. J. Lovas and C. Chu, 51st OSU Int. Symp. On Mol. Spectrosc., Columbus, OH, 1999, RB07.

bA. Vega-Toribio, A. Lesarri, R.D. Suenram, J. Grabow, 64th OSU Int. Symp. On Mol. Spectrosc., Columbus, OH, 2009, MH07.

cA. Lesarri, A. Vega-Toribio, R. D. Suenram, D. J. Brugh, J.-U. Grabow, Phys. Chem. Chem. Phys., 12, 9624-9631 (2010).